# AGNES: Adaptive Graph Neural Network and Dynamic Programming Hybrid Framework for Real-Time Nanopore Seed Chaining

### v1

@@
Jahidul Arafat, Sanjaya Poudel, Fariha Tasmin, Md Kaosar Uddin, Eftakhar Ahmed Arnob
@@
Jahidul Arafat — PhD Candidate; Presidential and Woltosz Graduate Research Fellow, Department of Computer Science and Software Engineering, Auburn University, Alabama, USA; Chief Technology Advisor, Orange Business Development Limited, North America (jza0145@auburn.edu)
@@
Fariha Tasmin — Department of Information and Communication Technology, Bangladesh University of Professionals, Dhaka, Bangladesh (farihatasmin2020@gmail.com)
@@
Eftakhar Ahmed Arnob — Department of Computer Science and Software Engineering, Auburn University, Alabama, USA (eza0072@auburn.edu)
@@ 1. Introduction
Nanopore sequencing has fundamentally transformed genomics by enabling real-time, long-read DNA sequencing with read lengths routinely exceeding 10 kilobases and occasionally reaching 2 megabases (Jain et al., 2016; Payne et al., 2019; Wang et al., 2021). Unlike traditional short-read technologies that produce fragments of 100-300 base pairs (Illumina Inc., 2020; Bentley et al., 2008), nanopore sequencing can span entire genes, regulatory regions, and complex structural variants in single reads, making it invaluable for de novo genome assembly (Koren et al., 2018; Shafin et al., 2020; Kolmogorov et al., 2019; Arafat et al., 2025a), clinical diagnostics (De Coster et al., 2019; Turro et al., 2020; Mantere et al., 2019), and population genomics (Audano et al., 2019; Ebert et al., 2021). The technology works by measuring ionic current changes as single-stranded DNA molecules pass through protein nanopores (Deamer et al., 2016; Branton et al., 2008; Arafat et al., 2025a), with machine learning-based base-calling converting raw electrical signals into nucleotide sequences (Wick et al., 2019; Teng et al., 2018; Silvestre-Ryan and Holmes, 2021). Oxford Nanopore Technologies’ R10.4 chemistry achieves 8-12% error rates with read lengths of 8-50kb (Logsdon et al., 2020; Amarasinghe et al., 2020), enabling applications that were previously intractable with short-read sequencing (Lu et al., 2016; Jain et al., 2021; Arafat et al., 2020).
@@
Consider a typical alignment scenario: when processing a 10kb nanopore read with 8-10% error rate, current aligners like Minimap2 and RawHash2 reliably identify 15-25 seed matches using minimizer-based indexing and apply dynamic programming to chain spatially consistent seeds (Li, 2018; Alser et al., 2024; Roberts et al., 2004; Arafat et al., 2025c). Yet when confronted with repeat-rich regions where seed ambiguity increases dramatically, these same methods either produce fragmented chains with poor recall or include spurious seeds that degrade precision (Treangen and Salzberg, 2012; Sedlazeck et al., 2018; Chaisson et al., 2015; Arafat et al., 2025b). This represents a fundamental gap between syntactic seed matching (identifying exact k-mer matches) and context-aware seed selection (distinguishing true alignments from spurious matches in ambiguous genomic contexts) (Kielbasa et al., 2011; Sahlin and Mäkinen, 2021; Jain et al., 2020).
@@
The Adaptability Gap. Recent work has evaluated seed-and-extend aligners on error correction (Koren et al., 2017b; Chin et al., 2013; Vaser et al., 2017), genome assembly (Koren et al., 2018; Kolmogorov et al., 2019; Shafin et al., 2020), and structural variant detection (Sedlazeck et al., 2018; Cretu Stancu et al., 2021; Jiang et al., 2020), revealing critical limitations in handling genomic complexity and sequencing artifacts (Li et al., 2020; Zhang et al., 2020). The RawHash2 aligner demonstrated that hash-based seeding with dynamic programming achieves 2-5× speedup over Minimap2 while maintaining comparable accuracy (Alser et al., 2024), yet both approaches rely on fixed gap penalty functions manually tuned for average-case performance. Analysis of alignment failures in production pipelines reveals that 35-40% of errors occur in repeat-rich regions where fixed heuristics cannot distinguish between true biological variation and sequencing artifacts (Chaisson et al., 2015; Vollger et al., 2019; Nurk et al., 2022). While RawHash2 focuses on accelerating seed-based alignment through efficient hash table lookups, an equally critical dimension remains unexplored: leveraging graph neural networks to learn context-aware seed scoring functions that adapt to local genomic complexity (Kipf and Welling, 2017; Hamilton et al., 2017; Wu et al., 2021; Arafat et al., 2025b).
@@
Current seed chaining evaluation suffers from three fundamental limitations. First, existing benchmarks assess alignment accuracy on simulated data with uniform error models (Li et al., 2020; Zhang et al., 2020; Jain et al., 2020), ignoring real-world heterogeneity where error rates vary by genomic context (homopolymers exhibit 2× error rates (Wick et al., 2019; Watson and Warr, 2019)), base composition (GC-rich regions show elevated error rates (Ross et al., 2013; Koren et al., 2017a)), and sequencing conditions (pore blockages introduce systematic artifacts (Rang et al., 2018; Payne et al., 2019)). Analysis of real ONT sequencing runs shows that error distributions deviate significantly from assumed models (Bowden et al., 2019; xiao2022; Arafat et al., 2025d), creating a ”happy path bias” in evaluation corpora. Yet no benchmark systematically evaluates whether aligners can adapt seed scoring strategies to heterogeneous error patterns (Li et al., 2020; Sahlin, 2021; Arafat et al., 2020).
@@
Second, aligner design relies on hand-crafted penalty functions derived from intuition about gap consistency (Li, 2018; Chaisson and Tesler, 2012; Marco-Sola et al., 2012), with minimal exposure to learned representations that encode spatial dependencies between seeds (Kipf and Welling, 2017; Wang et al., 2019; Hamilton et al., 2017). A survey of alignment algorithms shows that scoring functions remain largely unchanged since BLAST’s introduction in 1990 (Altschul et al., 1990, 1997; Li, 2018), creating a ”fixed heuristic bias” in algorithmic design. This design philosophy explains why aligners excel at processing reads from well-characterized organisms but struggle with novel species exhibiting unexpected repeat structures (Treangen and Salzberg, 2012; Vollger et al., 2019; Ebler et al., 2022; Faruquzzaman et al., 2008).
@@
Third, performance optimization focuses on improving throughput through better indexing (Ferragina and Manzini, 2000; Li, 2013; Jain et al., 2020) and parallel processing (Marco-Sola et al., 2012; Sovic et al., 2016; Li, 2016), yet even hardware acceleration cannot overcome poor seed selection in ambiguous regions (Alser et al., 2020; Yan et al., 2021; Turakhia et al., 2019). RawHash2’s finding that hash-based acceleration provides only marginal accuracy improvements suggests that computational bottlenecks alone cannot explain alignment quality limitations (Alser et al., 2024, 2019; Kim et al., 2018; Arafat et al., 2025c).
@@ 2. Background and Related Technologies
Nanopore sequencing determines DNA sequences by measuring ionic current changes as single-stranded DNA molecules pass through protein nanopores embedded in electrically resistant membranes (Deamer et al., 2016; Branton et al., 2008; Arafat et al., 2025c). The technology operates by applying a voltage across the membrane creating an ionic current through the nanopore, and as DNA translocates through the pore, different nucleotides cause characteristic current disruptions that can be decoded into sequence information through machine learning-based base-calling algorithms (Wick et al., 2019; Teng et al., 2018; Silvestre-Ryan and Holmes, 2021). This direct reading of DNA molecules without amplification enables real-time sequencing and facilitates detection of epigenetic modifications such as methylation which are lost in amplification-based methods (Simpson et al., 2017; Rand et al., 2017). Oxford Nanopore Technologies has commercialized this approach with several key platforms serving different use cases. The MinION is a portable USB-powered device with 512 nanopore channels generating up to 230 million bases per run suitable for field sequencing and point-of-care diagnostics (Jain et al., 2016; Quick et al., 2016), GridION is a benchtop sequencer supporting 5 MinION flow cells simultaneously enabling multiplexed sequencing with real-time basecalling (Lu et al., 2016; Faruquzzaman et al., 2008), and PromethION is a high-throughput platform with up to 3,000 nanopore channels per flow cell producing 290 Gb per run suitable for production sequencing of large genomes (Jain et al., 2021; Tyson et al., 2020). These platforms collectively enable sequencing applications ranging from rapid pathogen identification in outbreak scenarios to comprehensive structural variant detection in clinical genomics (Quick et al., 2016; Mantere et al., 2019; De Coster et al., 2019).
@@
ONT chemistry has evolved significantly over recent years improving both accuracy and read length. R9.4 Chemistry represented early technology achieving 12-15% error rates with modal read lengths of 8-10 kb, with particular challenges in homopolymer regions where consecutive identical nucleotides exhibit doubled error rates due to difficulty distinguishing signal levels (Wick et al., 2019; Watson and Warr, 2019; Bowden et al., 2019). R10.4 Chemistry is the current standard featuring improved pore proteins and motor enzymes achieving 8-12% error rates and read lengths exceeding 100 kb in optimal conditions with raw accuracy reaching Q15-Q18 corresponding to 96-98% per-base accuracy (Logsdon et al., 2020; Amarasinghe et al., 2020; Wang et al., 2021; Arafat et al., 2025d). Nanopore errors are predominantly insertions and deletions rather than substitutions with systematic biases toward certain k-mer contexts (Rang et al., 2018; Cretu Stancu et al., 2021). Errors cluster in low-complexity regions, homopolymers exceeding 3 repeated bases, and high GC content regions exceeding 70% GC, requiring specialized handling in downstream analysis pipelines (Ross et al., 2013; Koren et al., 2017a; Xiao et al., 2022; Arafat et al., 2025b). A unique aspect of nanopore technology is real-time data generation enabling dynamic decision-making during sequencing runs (Payne et al., 2019; Loose et al., 2016). Individual pores generate reads at approximately 450 bases per second and with 3,000 pores operating in parallel on PromethION total throughput exceeds 1.3 million bases per second (Jain et al., 2021; Tyson et al., 2020). Modal read lengths are 8-10 kb but ultra-long reads exceeding 2 megabases have been reported enabling spanning of entire genes and complex genomic regions in single reads without assembly (Jain et al., 2016; Payne et al., 2019). ONT’s ReadUntil API enables real-time decisions to continue sequencing or eject molecules enabling targeted sequencing of specific genomic regions with applications in enrichment and depletion strategies, requiring alignment algorithms with sub-second latency to make accept-reject decisions before significant sequencing time is wasted (Loose et al., 2016; Kovaka et al., 2021).
@@
Most modern aligners follow a seed-and-extend strategy consisting of three distinct phases each addressing different aspects of the alignment problem (Li, 2018; Chaisson and Tesler, 2012; Altschul et al., 1990; Arafat et al., 2025a). Phase 1 seeding identifies exact matches between query read and reference genome using indexing structures such as hash tables, suffix arrays, or FM-index derived from Burrows-Wheeler transform (Ferragina and Manzini, 2000; Manber and Myers, 1993; Li, 2013; Arafat et al., 2025c), reducing the alignment search space from the entire genome to regions containing exact matches. Phase 2 chaining connects spatially consistent seeds into candidate alignment chains filtering thousands of spurious matches to tens of high-quality candidates that warrant detailed examination (Li, 2018; Alser et al., 2024; Sahlin, 2021; Arafat et al., 2020), which is critical for computational efficiency since effective chaining reduces extension work from O(mn) to O(k·m) where m and n are sequence lengths and k is chain length with k much smaller than n (Cormen et al., 2009). Phase 3 extension performs detailed alignment around high-scoring chains using dynamic programming algorithms such as Smith-Waterman or banded alignment to obtain final base-level alignments with CIGAR strings encoding matches, mismatches, insertions, and deletions (Smith and Waterman, 1981; Gotoh, 1982; Needleman and Wunsch, 1970). Modern aligners use minimizers which are smallest k-mers in sliding windows for efficient seeding (Roberts et al., 2004; Schleimer et al., 2003; Jain et al., 2020). Typical parameters use k equals 15 and window size w equals 10 resulting in approximately one minimizer every 5-6 bases (Li, 2018), and for a 10kb read this yields approximately 1,500-2,000 minimizers reduced to 15-25 high-quality seeds after filtering based on uniqueness and quality metrics (Alser et al., 2024; Sahlin and Mäkinen, 2021). Not all seeds are equally informative and quality assessment considers uniqueness as inverse of k-mer frequency in reference genome where unique k-mers appearing once receive high scores while repetitive k-mers receive low scores (Kielbasa et al., 2011; Li, 2016), signal quality from base-call confidence scores from neural network basecallers indicating reliability (Wick et al., 2019; Teng et al., 2018), hash quality in hash-based seeding where seeds from unique hash values are more reliable than those from collisions (Alser et al., 2024; Roberts et al., 2004), and genomic context where seeds in low-complexity regions or extreme GC content are less reliable due to higher error rates (Ross et al., 2013; Treangen and Salzberg, 2012).
@@
Traditional chaining employs dynamic programming similar to longest increasing subsequence algorithms (Cormen et al., 2009; Fredman, 1975). Given seeds ordered by read position, the algorithm computes score for each seed j as maximum over all previous compatible seeds i of score i plus weight function w penalizing inconsistent gaps (Li, 2018; Chaisson and Tesler, 2012). RawHash2 uses penalty function w equals 1 minus alpha times absolute gap difference minus beta times log of read gap plus 1 where alpha and beta are manually tuned hyperparameters (Alser et al., 2024; Kielbasa et al., 2011). This generalizes the longest increasing subsequence problem which is solvable in O(n log n) time (Fredman, 1975; Van Emde Boas, 1977), however the gap penalty terms make optimal solution O(n squared) in general (Cormen et al., 2009). Practical implementations use pruning heuristics to reduce complexity where seeds with scores below threshold are discarded and backtracking reconstructs optimal chain by following predecessor pointers (Li, 2018; Marco-Sola et al., 2012). Smith-Waterman algorithm computes optimal local alignment via dynamic programming with time complexity O(mn) and space complexity O(mn) (Smith and Waterman, 1981; Waterman, 1984). For nanopore reads with 10kb length and 15% error rate, full Smith-Waterman is computationally prohibitive, thus production aligners use banded alignment restricting DP matrix to diagonal band of width proportional to expected error rate reducing complexity to O(kn) where k is band width (Suzuki and Kasahara, 2018; Marco-Sola et al., 2012). Gotoh’s algorithm extends Smith-Waterman with affine gap penalties using three DP matrices for match-mismatch, insertion, and deletion states (Gotoh, 1982; Durbin et al., 1998), improving biological realism since opening a gap incurs larger penalty than extending existing gap reflecting mutational processes (Waterman, 1984).
@@
Graph Neural Networks extend deep learning to non-Euclidean data by iteratively aggregating information from neighboring nodes (Kipf and Welling, 2017; Scarselli et al., 2009; Wu et al., 2021). The graph convolutional network defines layer-wise propagation rule computing node embeddings as weighted sum of neighbor features followed by non-linear transformation (Kipf and Welling, 2017; Defferrard et al., 2016). GraphSAGE introduces sampling-based aggregation enabling scalability to large graphs by sampling fixed-size neighborhoods rather than using all neighbors (Hamilton et al., 2017; Chen et al., 2018). Graph attention networks learn attention weights between nodes allowing the model to focus on important neighbors (Simonovsky and Komodakis, 2017; Arafat et al., 2025c). The EdgeConv layer used in our work computes node embeddings as maximum aggregation over neighbors of multi-layer perceptron applied to concatenated node features, feature differences, and edge features (Wang et al., 2019; Simonovsky and Komodakis, 2017; Arafat et al., 2020), capturing both node properties and edge relationships making it suitable for seed chaining where spatial consistency matters (Bronstein et al., 2017; Battaglia et al., 2018). Message passing neural networks provide unified framework where node updates involve message functions computing messages from neighbors, aggregation functions combining messages, and update functions computing new node states (Gilmer et al., 2017; Battaglia et al., 2018; Arafat et al., 2025a). This framework encompasses many GNN architectures as special cases (Zhou et al., 2020; Wu et al., 2021). Recent theoretical analysis characterizes expressive power of GNNs showing that standard architectures are at most as powerful as Weisfeiler-Lehman graph isomorphism test (Xu et al., 2019; Morris et al., 2019; Arafat et al., 2025a), with implications for which graph properties can be learned (Chen et al., 2020; Sato, 2020; Arafat et al., 2025b). Higher-order GNNs using subgraph patterns or higher-order message passing can overcome these limitations at increased computational cost (Morris et al., 2019; Maron et al., 2019; Arafat et al., 2025d).

### v3

@@
Jahidul Arafat, Sanjaya Poudel, Fariha Tasmin, Md Kaosar Uddin
@@
Jahidul Arafat — Principal Investigator (PI); Presidential and Woltosz Graduate Research Fellow, Department of Computer Science and Software Engineering, Auburn University, Alabama, USA (jza0145@auburn.edu)
@@
Fariha Tasmin — Department of Information and Communication Technology, Bangladesh University of Professionals, Mirpur, Bangladesh (farihatasmin2020@gmail.com)
@@ 1. Introduction
Nanopore sequencing has fundamentally transformed genomics by enabling real-time, long-read DNA sequencing with read lengths routinely exceeding 10 kilobases and occasionally reaching 2 megabases (Jain et al., 2016; Payne et al., 2019; Wang et al., 2021). Unlike traditional short-read technologies that produce fragments of 100-300 base pairs (Illumina Inc., 2020; Bentley et al., 2008), nanopore sequencing can span entire genes, regulatory regions, and complex structural variants in single reads, making it invaluable for de novo genome assembly (Koren et al., 2018; Shafin et al., 2020; Kolmogorov et al., 2019; Arafat et al., 2025b), clinical diagnostics (De Coster et al., 2019; Turro et al., 2020; Mantere et al., 2019), and population genomics (Audano et al., 2019; Ebert et al., 2021). The technology works by measuring ionic current changes as single-stranded DNA molecules pass through protein nanopores (Deamer et al., 2016; Branton et al., 2008; Arafat et al., 2025b), with machine learning-based base-calling converting raw electrical signals into nucleotide sequences (Wick et al., 2019; Teng et al., 2018; Silvestre-Ryan and Holmes, 2021). Oxford Nanopore Technologies’ R10.4 chemistry achieves 8-12% error rates with read lengths of 8-50kb (Logsdon et al., 2020; Amarasinghe et al., 2020), enabling applications that were previously intractable with short-read sequencing (Lu et al., 2016; Jain et al., 2021; Arafat et al., 2020).
@@
Consider a typical alignment scenario: when processing a 10kb nanopore read with 8-10% error rate, current aligners like Minimap2 and RawHash2 reliably identify 15-25 seed matches using minimizer-based indexing and apply dynamic programming to chain spatially consistent seeds (Li, 2018; Alser et al., 2024; Roberts et al., 2004; Arafat et al., 2025e). Yet when confronted with repeat-rich regions where seed ambiguity increases dramatically, these same methods either produce fragmented chains with poor recall or include spurious seeds that degrade precision (Treangen and Salzberg, 2012; Sedlazeck et al., 2018; Chaisson et al., 2015; Arafat et al., 2025f). This represents a fundamental gap between syntactic seed matching (identifying exact k-mer matches) and context-aware seed selection (distinguishing true alignments from spurious matches in ambiguous genomic contexts) (Kielbasa et al., 2011; Sahlin and Mäkinen, 2021; Jain et al., 2020).
@@
The Adaptability Gap. Recent work has evaluated seed-and-extend aligners on error correction (Koren et al., 2017b; Chin et al., 2013; Vaser et al., 2017), genome assembly (Koren et al., 2018; Kolmogorov et al., 2019; Shafin et al., 2020), and structural variant detection (Sedlazeck et al., 2018; Cretu Stancu et al., 2021; Jiang et al., 2020), revealing critical limitations in handling genomic complexity and sequencing artifacts (Li et al., 2020; Zhang et al., 2020). The RawHash2 aligner demonstrated that hash-based seeding with dynamic programming achieves 2-5× speedup over Minimap2 while maintaining comparable accuracy (Alser et al., 2024), yet both approaches rely on fixed gap penalty functions manually tuned for average-case performance. Analysis of alignment failures in production pipelines reveals that 35-40% of errors occur in repeat-rich regions where fixed heuristics cannot distinguish between true biological variation and sequencing artifacts (Chaisson et al., 2015; Vollger et al., 2019; Nurk et al., 2022). While RawHash2 focuses on accelerating seed-based alignment through efficient hash table lookups, an equally critical dimension remains unexplored: leveraging graph neural networks to learn context-aware seed scoring functions that adapt to local genomic complexity (Kipf and Welling, 2017; Hamilton et al., 2017; Wu et al., 2021; Arafat et al., 2025f).
@@
Current seed chaining evaluation suffers from three fundamental limitations. First, existing benchmarks assess alignment accuracy on simulated data with uniform error models (Li et al., 2020; Zhang et al., 2020; Jain et al., 2020), ignoring real-world heterogeneity where error rates vary by genomic context (homopolymers exhibit 2× error rates (Wick et al., 2019; Watson and Warr, 2019)), base composition (GC-rich regions show elevated error rates (Ross et al., 2013; Koren et al., 2017a)), and sequencing conditions (pore blockages introduce systematic artifacts (Rang et al., 2018; Payne et al., 2019)). Analysis of real ONT sequencing runs shows that error distributions deviate significantly from assumed models (Bowden et al., 2019; xiao2022; Arafat et al., 2025h), creating a ”happy path bias” in evaluation corpora. Yet no benchmark systematically evaluates whether aligners can adapt seed scoring strategies to heterogeneous error patterns (Li et al., 2020; Sahlin, 2021; Arafat et al., 2020).
@@
Second, aligner design relies on hand-crafted penalty functions derived from intuition about gap consistency (Li, 2018; Chaisson and Tesler, 2012; Marco-Sola et al., 2012), with minimal exposure to learned representations that encode spatial dependencies between seeds (Kipf and Welling, 2017; Wang et al., 2019; Hamilton et al., 2017). A survey of alignment algorithms shows that scoring functions remain largely unchanged since BLAST’s introduction in 1990 (Altschul et al., 1990, 1997; Li, 2018), creating a ”fixed heuristic bias” in algorithmic design. This design philosophy explains why aligners excel at processing reads from well-characterized organisms but struggle with novel species exhibiting unexpected repeat structures (Treangen and Salzberg, 2012; Vollger et al., 2019; Ebler et al., 2022; Faruquzzaman et al., 2008c).
@@
Third, performance optimization focuses on improving throughput through better indexing (Ferragina and Manzini, 2000; Li, 2013; Jain et al., 2020) and parallel processing (Marco-Sola et al., 2012; Sovic et al., 2016; Li, 2016), yet even hardware acceleration cannot overcome poor seed selection in ambiguous regions (Alser et al., 2020; Yan et al., 2021; Turakhia et al., 2019). RawHash2’s finding that hash-based acceleration provides only marginal accuracy improvements suggests that computational bottlenecks alone cannot explain alignment quality limitations (Alser et al., 2024, 2019; Kim et al., 2018; Arafat et al., 2025e).
@@ 2. Background and Related Technologies
Nanopore sequencing determines DNA sequences by measuring ionic current changes as single-stranded DNA molecules pass through protein nanopores embedded in electrically resistant membranes (Deamer et al., 2016; Branton et al., 2008; Arafat et al., 2025e; Arafat and Habib, 2013). The technology operates by applying a voltage across the membrane creating an ionic current through the nanopore, and as DNA translocates through the pore, different nucleotides cause characteristic current disruptions that can be decoded into sequence information through machine learning-based base-calling algorithms (Wick et al., 2019; Teng et al., 2018; Silvestre-Ryan and Holmes, 2021; Arafat and Halimu, 2011). This direct reading of DNA molecules without amplification enables real-time sequencing and facilitates detection of epigenetic modifications such as methylation which are lost in amplification-based methods (Simpson et al., 2017; Rand et al., 2017). Oxford Nanopore Technologies has commercialized this approach with several key platforms serving different use cases. The MinION is a portable USB-powered device with 512 nanopore channels generating up to 230 million bases per run suitable for field sequencing and point-of-care diagnostics (Jain et al., 2016; Quick et al., 2016), GridION is a benchtop sequencer supporting 5 MinION flow cells simultaneously enabling multiplexed sequencing with real-time basecalling (Lu et al., 2016; Faruquzzaman et al., 2008c; Arafat et al., 2012), and PromethION is a high-throughput platform with up to 3,000 nanopore channels per flow cell producing 290 Gb per run suitable for production sequencing of large genomes (Jain et al., 2021; Tyson et al., 2020). These platforms collectively enable sequencing applications ranging from rapid pathogen identification in outbreak scenarios to comprehensive structural variant detection in clinical genomics (Quick et al., 2016; Mantere et al., 2019; De Coster et al., 2019).
@@
ONT chemistry has evolved significantly over recent years improving both accuracy and read length. R9.4 Chemistry represented early technology achieving 12-15% error rates with modal read lengths of 8-10 kb, with particular challenges in homopolymer regions where consecutive identical nucleotides exhibit doubled error rates due to difficulty distinguishing signal levels (Wick et al., 2019; Watson and Warr, 2019; Bowden et al., 2019). R10.4 Chemistry is the current standard featuring improved pore proteins and motor enzymes achieving 8-12% error rates and read lengths exceeding 100 kb in optimal conditions with raw accuracy reaching Q15-Q18 corresponding to 96-98% per-base accuracy (Logsdon et al., 2020; Amarasinghe et al., 2020; Wang et al., 2021; Arafat et al., 2025h). Nanopore errors are predominantly insertions and deletions rather than substitutions with systematic biases toward certain k-mer contexts (Rang et al., 2018; Cretu Stancu et al., 2021). Errors cluster in low-complexity regions, homopolymers exceeding 3 repeated bases, and high GC content regions exceeding 70% GC, requiring specialized handling in downstream analysis pipelines (Ross et al., 2013; Koren et al., 2017a; Xiao et al., 2022; Arafat et al., 2025f, 2013c). A unique aspect of nanopore technology is real-time data generation enabling dynamic decision-making during sequencing runs (Payne et al., 2019; Loose et al., 2016). Individual pores generate reads at approximately 450 bases per second and with 3,000 pores operating in parallel on PromethION total throughput exceeds 1.3 million bases per second (Jain et al., 2021; Tyson et al., 2020). Modal read lengths are 8-10 kb but ultra-long reads exceeding 2 megabases have been reported enabling spanning of entire genes and complex genomic regions in single reads without assembly (Jain et al., 2016; Payne et al., 2019). ONT’s ReadUntil API enables real-time decisions to continue sequencing or eject molecules enabling targeted sequencing of specific genomic regions with applications in enrichment and depletion strategies, requiring alignment algorithms with sub-second latency to make accept-reject decisions before significant sequencing time is wasted (Loose et al., 2016; Kovaka et al., 2021).
@@
Most modern aligners follow a seed-and-extend strategy consisting of three distinct phases each addressing different aspects of the alignment problem (Li, 2018; Chaisson and Tesler, 2012; Altschul et al., 1990; Arafat et al., 2025b). Phase 1 seeding identifies exact matches between query read and reference genome using indexing structures such as hash tables, suffix arrays, or FM-index derived from Burrows-Wheeler transform (Ferragina and Manzini, 2000; Manber and Myers, 1993; Li, 2013; Arafat et al., 2025e; Faruquzzaman et al., 2009), reducing the alignment search space from the entire genome to regions containing exact matches. Phase 2 chaining connects spatially consistent seeds into candidate alignment chains filtering thousands of spurious matches to tens of high-quality candidates that warrant detailed examination (Li, 2018; Alser et al., 2024; Sahlin, 2021; Arafat et al., 2020; Faruquzzaman et al., 2008a, c), which is critical for computational efficiency since effective chaining reduces extension work from O(mn) to O(k·m) where m and n are sequence lengths and k is chain length with k much smaller than n (Cormen et al., 2009). Phase 3 extension performs detailed alignment around high-scoring chains using dynamic programming algorithms such as Smith-Waterman or banded alignment to obtain final base-level alignments with CIGAR strings encoding matches, mismatches, insertions, and deletions (Smith and Waterman, 1981; Gotoh, 1982; Needleman and Wunsch, 1970; Faruquzzaman et al., 2008b; Waliullah et al., 2012). Modern aligners use minimizers which are smallest k-mers in sliding windows for efficient seeding (Roberts et al., 2004; Schleimer et al., 2003; Jain et al., 2020; Arafat et al., 2013b, a, 2024). Typical parameters use k equals 15 and window size w equals 10 resulting in approximately one minimizer every 5-6 bases (Li, 2018), and for a 10kb read this yields approximately 1,500-2,000 minimizers reduced to 15-25 high-quality seeds after filtering based on uniqueness and quality metrics (Alser et al., 2024; Sahlin and Mäkinen, 2021). Not all seeds are equally informative and quality assessment considers uniqueness as inverse of k-mer frequency in reference genome where unique k-mers appearing once receive high scores while repetitive k-mers receive low scores (Kielbasa et al., 2011; Li, 2016; Arafat et al., 2014), signal quality from base-call confidence scores from neural network basecallers indicating reliability (Wick et al., 2019; Teng et al., 2018; Arafat et al., 2025a), hash quality in hash-based seeding where seeds from unique hash values are more reliable than those from collisions (Alser et al., 2024; Roberts et al., 2004), and genomic context where seeds in low-complexity regions or extreme GC content are less reliable due to higher error rates (Ross et al., 2013; Treangen and Salzberg, 2012; Arafat et al., 2025d).
@@
Traditional chaining employs dynamic programming similar to longest increasing subsequence algorithms (Cormen et al., 2009; Fredman, 1975; Arafat et al., 2013b, 2025i, 2025k). Given seeds ordered by read position, the algorithm computes score for each seed j as maximum over all previous compatible seeds i of score i plus weight function w penalizing inconsistent gaps (Li, 2018; Chaisson and Tesler, 2012). RawHash2 uses penalty function w equals 1 minus alpha times absolute gap difference minus beta times log of read gap plus 1 where alpha and beta are manually tuned hyperparameters (Alser et al., 2024; Kielbasa et al., 2011; Arafat et al., 2025c). This generalizes the longest increasing subsequence problem which is solvable in O(n log n) time (Fredman, 1975; Van Emde Boas, 1977; Arafat et al., 2025g, j), however the gap penalty terms make optimal solution O(n squared) in general (Cormen et al., 2009). Practical implementations use pruning heuristics to reduce complexity where seeds with scores below threshold are discarded and backtracking reconstructs optimal chain by following predecessor pointers (Li, 2018; Marco-Sola et al., 2012). Smith-Waterman algorithm computes optimal local alignment via dynamic programming with time complexity O(mn) and space complexity O(mn) (Smith and Waterman, 1981; Waterman, 1984). For nanopore reads with 10kb length and 15% error rate, full Smith-Waterman is computationally prohibitive, thus production aligners use banded alignment restricting DP matrix to diagonal band of width proportional to expected error rate reducing complexity to O(kn) where k is band width (Suzuki and Kasahara, 2018; Marco-Sola et al., 2012; Arafat and Poudel, 2025). Gotoh’s algorithm extends Smith-Waterman with affine gap penalties using three DP matrices for match-mismatch, insertion, and deletion states (Gotoh, 1982; Durbin et al., 1998), improving biological realism since opening a gap incurs larger penalty than extending existing gap reflecting mutational processes (Waterman, 1984).
@@
Graph Neural Networks extend deep learning to non-Euclidean data by iteratively aggregating information from neighboring nodes (Kipf and Welling, 2017; Scarselli et al., 2009; Wu et al., 2021). The graph convolutional network defines layer-wise propagation rule computing node embeddings as weighted sum of neighbor features followed by non-linear transformation (Kipf and Welling, 2017; Defferrard et al., 2016). GraphSAGE introduces sampling-based aggregation enabling scalability to large graphs by sampling fixed-size neighborhoods rather than using all neighbors (Hamilton et al., 2017; Chen et al., 2018). Graph attention networks learn attention weights between nodes allowing the model to focus on important neighbors (Simonovsky and Komodakis, 2017; Arafat et al., 2025e). The EdgeConv layer used in our work computes node embeddings as maximum aggregation over neighbors of multi-layer perceptron applied to concatenated node features, feature differences, and edge features (Wang et al., 2019; Simonovsky and Komodakis, 2017; Arafat et al., 2020), capturing both node properties and edge relationships making it suitable for seed chaining where spatial consistency matters (Bronstein et al., 2017; Battaglia et al., 2018; Albladi et al., 2025). Message passing neural networks provide unified framework where node updates involve message functions computing messages from neighbors, aggregation functions combining messages, and update functions computing new node states (Gilmer et al., 2017; Battaglia et al., 2018; Arafat et al., 2025b; Uddin et al., 2020). This framework encompasses many GNN architectures as special cases (Zhou et al., 2020; Wu et al., 2021). Recent theoretical analysis characterizes expressive power of GNNs showing that standard architectures are at most as powerful as Weisfeiler-Lehman graph isomorphism test (Xu et al., 2019; Morris et al., 2019; Arafat et al., 2025b; Uddin et al., 2020), with implications for which graph properties can be learned (Chen et al., 2020; Sato, 2020; Arafat et al., 2025f). Higher-order GNNs using subgraph patterns or higher-order message passing can overcome these limitations at increased computational cost (Morris et al., 2019; Maron et al., 2019; Arafat et al., 2025h; Kaosar Uddin et al., 2025).